# Influence-aware Causal Autoencoder Network for Node Importance Ranking in Complex Networks

### v1

@@ Abstract
Node importance ranking is a fundamental problem in graph data analysis. Existing approaches typically rely on node features derived from either traditional centrality measures or advanced graph representation learning methods, which depend directly on the target network’s topology. However, this reliance on structural information raises privacy concerns and often leads to poor generalization across different networks. In this work, we address a key question: Can we design a node importance ranking model trained exclusively on synthetic networks that is effectively appliable to real-world networks, eliminating the need to rely on the topology of target networks and improving both practicality and generalizability? We answer this question affirmatively by proposing the Influence-aware Causal Autoencoder Network (ICAN), a novel framework that leverages causal representation learning to get robust, invariant node embeddings for cross-network ranking tasks. Firstly, ICAN introduces an influence-aware causal representation learning module within an autoencoder architecture to extract node embeddings that are causally related to node importance. Moreover, we introduce a causal ranking loss and design a unified optimization framework that jointly optimizes the reconstruction and ranking objectives, enabling mutual reinforcement between node representation learning and ranking optimization. This design allows ICAN, trained on synthetic networks, to generalize effectively across diverse real-world graphs. Extensive experiments on multiple benchmark datasets demonstrate that ICAN consistently outperforms state-of-the-art baselines in terms of both ranking accuracy and generalization capability.
I Introduction
@@
Critical nodes in complex networks are of fundamental importance, as they play key roles in maintaining the network’s functionality, performance, stability, robustness, and dynamic behavior [1]. For example, in social networks, they facilitate targeted information dissemination and optimized resource allocation [2]. In transportation networks, they enable vulnerability analysis and robustness enhancement [3]. In biological networks, they help identify essential genes or proteins for therapeutic interventions [4]. With the growing scale and complexity of real-world networks, accurately identifying critical nodes has become an increasingly important yet challenging task [5, 6, 7, 8]. Developing methods that are both accurate and generalizable for node importance ranking is therefore of great theoretical and practical value [9, 10].
@@
There has been significant research on node importance ranking in complex networks. Traditional topology-based approaches primarily rely on centrality-based measures, such as degree centrality [11], eigenvector centrality [12], and betweenness centrality [12, 13]. While some centrality measures are straightforward to compute (e.g., degree centrality), others (e.g., betweenness) can be computationally demanding, particularly in large-scale networks. Moreover, these methods focus on computing individual importance scores rather than modeling relative importance relations among nodes, which limits their effectiveness in ranking tasks. In addition, most measures capture node importance along a single structural dimension and are often tailored to specific research goals, which limits their generalizability across diverse network topologies. For instance, the degree-centrality-based method can identify high-degree hubs as the most influential nodes in a network. However, such an assumption overlooks the fact that influence can arise from nodes occupying structurally balanced or strategically positioned roles, even when their degrees are moderate.
@@
In recent years, deep representation learning–based methods have become a powerful paradigm for node ranking. By automatically learning expressive node embeddings, these methods enable downstream tasks such as regression to be efficiently performed on the learned embeddings. This category includes Graph Convolutional Networks (GCNs) for local neighborhood aggregation [14], graph embedding techniques for dimensionality reduction [15, 16], and more sophisticated architectures such as graph attention networks [17] and graph contrastive learning frameworks [18] that capture nuanced structural dependencies.
@@
Despite their strong representational power, most existing deep representation learning methods focus solely on modeling network structures—capturing low- and high-order proximities between nodes—while overlooking node importance information that is crucial for ranking tasks. For example, AGNN [16] and CGNN [19] typically follow a two-stage paradigm, where node representations are first learned independently using graph neural network techniques, followed by a prediction module for ranking or regression. As task-specific objectives are not integrated into the representation learning stage, the resulting embeddings often fail to capture task-relevant information, leading to suboptimal performance in node importance ranking.
@@
Moreover, these deep representation learning–based approaches typically rely heavily on the topology of the target network. However, in many real-world scenarios, privacy constraints render the network a black box to users, who have no access to its complete structure [20]. Learning node embeddings directly from the explicit graph structure also limits the model’s ability to generalize across different networks, as such representations often overfit to specific structural patterns. These challenges motivate the development of node embedding methods that can improve the performance of importance ranking models without relying on the topology of target networks. In this work, we specifically investigate whether such embeddings can be learned through representation learning trained exclusively on synthetic networks and then applied across multiple diverse real-world networks.
@@
Objectives and challenges. To address the limitations of existing approaches, we propose a novel representation learning framework for node importance ranking in complex networks. Our method draws inspiration from causal representation learning [21, 22, 23], which provides domain-invariant representations that generalize across different environments. As a result, it can be trained on synthetic networks and generalize effectively to diverse real-world graphs. Specifically, we pursue two key objectives: (i) to obtain effective node representation suitable for ranking without using the target network structure—enabling applicability in practical scenarios; and (ii) to improve ranking performance through mutual reinforcement between node representation learning and ranking optimization—ensuring more reliable and interpretable ranking results. Achieving these goals is nontrivial, as it raises two major challenges: (i) how to design an effective unsupervised representation learning strategy that produces network-invariant node embeddings, and (ii) how to ensure that the learned representations are inherently relevant and informative for node ranking.
@@
Our solution. To address the first challenge, we design an influence-aware causal structure learning module within the autoencoder to capture robust low-dimensional embeddings that exhibit causal relevance to node importance. This mechanism enables the learned representations to be network-invariant and to generalize effectively to unseen target graphs. To tackle the second challenge, we formulate a unified objective function that jointly optimizes our proposed causal reconstruction and causal ranking losses, enabling a synergistic interaction between representation learning and ranking optimization.
Contributions. Our contributions are summarized as follows:
@@
  • Influence-aware causal representation learning mechanism: We design an influence-aware causal representation learning mechanism for node ranking prediction and integrate it into an autoencoder framework. Specifically, we construct a node influence variable based on the information propagation process within the training network, and learn the causal relationships between node embeddings and this influence variable. This design encourages the learned representations to capture network-invariant causal signals related to node importance, enabling the model to be trained solely on synthetic networks and to generalize effectively to diverse real-world graphs.
@@
  • Feature–task co-optimization mechanism: We propose the causal reconstruction loss and the causal ranking loss, and integrate them with a regularization term into a unified objective. This feature-task co-optimization framework jointly optimizes node representation learning and ranking prediction, ensuring that the resulting embeddings are directly aligned with the downstream ranking task.
@@
The remainder of this paper is organized as follows. Problem formulation is introduced in Section II. The proposed influence-aware causal autoencoder network for node importance ranking is presented in detail in Section III. The experimental results are reported in Section IV. Some related work is outlined in Section V. Conclusions are drawn in the final section.
II Problem formulation
@@
In this section, we define the problem of node ranking in a graph and introduce the key assumptions of the proposed model.
II-A Problem statement
@@
Given a graph $G=\{V,E\}$, where $V$ is the set of nodes and $E$ is the set of edges. The notation $v_{i}\in V$ denotes a node in $V$, and $e_{ij}\in E$ denotes an edge from node $v_{i}$ to $v_{j}$. The graph can be represented by an $n\times n$ adjacency matrix $\bm{A}$, where $n$ denotes the number of nodes. The entries of the matrix are defined such that $\bm{A}_{ij}=1$ when $e_{ij}\in E$, and $\bm{A}_{ij}=0$ otherwise. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix. For graphs without node attributes, $\bm{X}$ can be obtained by applying a node embedding method (e.g., Node2Vec [24]) to the adjacency matrix $\bm{A}$.
@@
The node importance ranking is a critical problem in network analysis. Given the graph $G=\{V,E\}$, our task is to assign a score $s_{i}$ to each node $v_{i}\in V$, forming a score vector $\bm{S}=(s_{1},s_{2},\dots,s_{n})$, These scores are then used to derive a node importance ranking $\bm{R}=(r_{1},r_{2},\dots,r_{n})$, where $r_{i}$ denotes the rank of node $v_{i}$. The ranking $R$ satisfies the following: $r_{i}=r_{j}$ if $s_{i}=s_{j}$, and $r_{i}<r_{j}$ if $s_{i}<s_{j}$, meaning that a higher score corresponds to a higher (i.e., numerically lower) rank. Unlike methods that emphasize the absolute magnitude of scores, our approach focuses on preserving the relative ordering of nodes consistent with their underlying importance, making ranking consistency the primary objective.
II-B Assumptions

### v2

# Importance Ranking in Complex Networks via Influence-aware Causal Node Embedding

@@ Abstract
Understanding and quantifying node importance is a fundamental problem in network science and engineering, underpinning a wide range of applications such as influence maximization, social recommendation, and network dismantling. Prior research often relies on centrality measures or advanced graph embedding techniques using structural information, followed by downstream classification or regression tasks to identify critical nodes. However, these methods typically decouple node representation learning from the ranking objective and rely on the topological structure of target networks, leading to feature–task inconsistency and limited generalization across networks. This paper proposes a novel framework that leverages causal representation learning to get robust, invariant node embeddings for cross-network ranking tasks. Firstly, we introduce an influence-aware causal node embedding module within an autoencoder architecture to extract node embeddings that are causally related to node importance. Moreover, we introduce a causal ranking loss and design a unified optimization framework that jointly optimizes the reconstruction and ranking objectives, enabling mutual reinforcement between node representation learning and ranking optimization. This design allows the proposed model to be trained on synthetic networks and to generalize effectively across diverse real-world networks. Extensive experiments on multiple benchmark datasets demonstrate that the proposed model consistently outperforms state-of-the-art baselines in terms of both ranking accuracy and cross-network transferability, offering new insights for network analysis and engineering applications—particularly in scenarios where the target network’s structure is inaccessible in advance due to privacy or security constraints.
Index Terms:
Complex networks, node importance, graph representation learning, node ranking, causal representation learning
1 Introduction
@@
Complex networks provide a powerful framework for modeling and analyzing a wide range of systems across diverse domains, including social networks, transportation systems, and biological networks [costa2011analyzing]. In these networks, nodes represent entities within a real system such as individuals, infrastructure components, or functional units, while edges capture interactions or relationships between them. A key challenge in network science and engineering is identifying important nodes, as they play pivotal roles in maintaining network functionality, performance, stability, and robustness [kaili2025finding]. For example, in social networks, they facilitate targeted information dissemination and optimized resource allocation [zhang2022influence]. In transportation networks, they enable vulnerability analysis and robustness enhancement [zhou2022analyzing]. In biological networks, they help identify essential genes or proteins for therapeutic interventions [de2023more]. With the growing scale and complexity of real-world networks, accurately identifying critical nodes has become an increasingly important yet challenging task [ma2025node]. Hence, developing accurate and generalizable methods for node importance ranking holds substantial theoretical and practical significance [liu2023identify].
@@
There has been significant research on node importance ranking in complex networks. Traditional topology-based approaches primarily rely on centrality-based measures, such as degree centrality [bavelas1948mathematical], eigenvector centrality [bonacich1972factoring], and betweenness centrality [xu2025adaptive]. While some centrality measures are straightforward to compute (e.g., degree centrality), others (e.g., betweenness) can be computationally demanding, particularly in large-scale networks. Moreover, these methods focus on computing individual importance scores rather than modeling relative importance relations among nodes, which limits their effectiveness in ranking tasks. In addition, most measures capture node importance along a single structural dimension and are often tailored to specific research goals, which limits their generalizability across diverse network topologies. For instance, the degree-centrality-based method can identify high-degree hubs as the most influential nodes in a network. However, such an assumption overlooks the fact that influence can arise from nodes occupying structurally balanced or strategically positioned roles, even when their degrees are moderate.
@@
In recent years, deep representation learning–based methods have become a powerful paradigm for node ranking. By automatically learning expressive node embeddings, these methods enable downstream tasks such as regression to be efficiently performed on the learned embeddings. This category includes Graph Convolutional Networks (GCNs) for local neighborhood aggregation [zhao2020infgcn], graph embedding techniques for dimensionality reduction [zhu2021community], and more sophisticated architectures such as graph attention networks [park2019estimating] and graph contrastive learning frameworks [liu2023node] that capture nuanced structural dependencies.
@@
Despite their strong representational power, most existing deep representation learning methods focus solely on modeling network structures—capturing low- and high-order proximities between nodes—while overlooking node importance information that is crucial for ranking tasks. For example, AGNN [xiong2024vital] and CGNN [zhang2022new] typically follow a two-stage paradigm, where node representations are first learned independently using graph neural network techniques, followed by a prediction module for ranking or regression. As task-specific objectives are not integrated into the representation learning stage, the resulting embeddings often fail to capture task-relevant information, leading to suboptimal performance in node importance ranking.
@@
Moreover, these deep representation learning–based approaches typically rely heavily on the topology of the target network. However, in many real-world scenarios, privacy or security constraints make the network effectively a black box to users, restricting access to its underlying structure in advance [li2021manipulating]. For instance, in various strategic or security-sensitive contexts, identifying critical nodes within a non-cooperative network is essential for understanding or mitigating its functionality, yet the network’s topology is typically unknown in advance due to intelligence and security limitations. Such situations exemplify a generic problem in network science and engineering—how to identify and rank critical nodes when the topology of the target network is unobservable. In this work, we specifically investigate whether feature embeddings that capture node importance can be learned through a representation learning model trained exclusively on synthetic networks and then transferred across multiple diverse real-world networks.
@@
To achieve this goal, our approach is inspired from causal representation learning [chu2023continual, zhou2025learning], which provides domain-invariant representations that generalize across different environments. Specifically, we pursue two key objectives: (i) to obtain effective node representation suitable for ranking without using the target network structure—enabling applicability in practical scenarios; and (ii) to improve ranking performance through mutual reinforcement between node representation learning and ranking optimization—ensuring more reliable and interpretable ranking results. Achieving these goals is nontrivial, as it raises two major challenges: (i) how to design an effective unsupervised representation learning strategy that produces network-invariant node embeddings, and (ii) how to ensure that the learned representations are inherently relevant and informative for node ranking.
@@
To address the first challenge, we design an influence-aware causal node embedding module within the autoencoder to capture robust low-dimensional embeddings that exhibit causal relevance to node importance. This mechanism enables the learned representations to be network-invariant and to generalize effectively to unseen target graphs. To tackle the second challenge, we formulate a unified objective function that jointly optimizes our proposed causal reconstruction and causal ranking losses, enabling a synergistic interaction between representation learning and ranking optimization. Our contributions can be summarized as follows:
@@
  • Influence-aware causal node embedding mechanism: We design an influence-aware causal node embedding mechanism for node ranking prediction and integrate it into an autoencoder framework. Specifically, we construct a node influence variable based on the information propagation process within the training network, and learn the causal relationships between node embeddings and this influence variable. This design encourages the learned node representations to capture network-invariant causal signals related to node importance, enabling the model to be trained solely on synthetic networks and to generalize effectively to diverse real-world graphs.
@@
  • Feature–task co-optimization mechanism: We propose a causal reconstruction loss and a causal ranking loss, and integrate them with a regularization term into a unified objective. This feature-task co-optimization framework jointly optimizes node representation learning and ranking prediction, ensuring that the resulting embeddings are directly aligned with the downstream ranking task.
@@
The remainder of this paper is organized as follows. Some related work is outlined in Section 2. Problem formulation is introduced in Section 3. The proposed influence-aware causal autoencoder model for node importance ranking is presented in detail in Section 4. The experimental results are reported in Section 5. Conclusions are drawn in the final section.
2 Related work
TABLE I: Comparative analysis of different node importance ranking methods.
Categories Methods Contributions Limitations
Neighborhood-based DC [nieminen1974centrality], $H$-index [hirsch2005index], $K$-shell [kitsak2010identification] Focus on local structural information. These methods measure node importance from a single structural dimension, which limits their applicability to more complex network structures.
Eigenvector-based EC [bonacich1987power], PageRank [page1998pagerank] Leverage spectral characteristics and iterative propagation mechanisms to evaluate global node influence.
Path-based BC [freeman1977set], ECC [hage1995eccentricity], CC [freeman1978centrality], ViralRank [iannelli2018influencers] Assess node importance via paths, distances, or random walks to capture multi-scale topological information.
Fusion-based ISCL [mohammad2024convex], HCT [wang2024identification], CBGN [liu2023influence], ITFRSGC [an2025identifying] Blend multiple centrality metrics or topological characteristics to create hybrid measures. These methods have high computational complexity and are sensitive to parameter tuning, demanding manual adjustment.
Machine learning-based ML [hajarathaiah2024node], LS-SVM [wen2018fast], MFIM [wang2024multi] Identify influential nodes by formulating the task as a regression or classification problem. These methods rely on explicit network topology information, limiting their generalization performance across networks.
Deep learning-based InfGCN [zhao2020infgcn], NDM [ahmad2024neural], CGNN [zhang2022new], ReGCN [wu2024graph] Integrate handcrafted features with deep learning architectures for influential node identification. These methods rely on handcrafted features and are unable to capture complex structural patterns.
AGNN [xiong2024vital], GCRA [kaili2025finding], MVPL [ma2025node], RCNN [yu2020identifying], STGCN [huang2025sparse] Directly generate latent node embeddings from graph topology and subsequently calculate importance scores. These methods focus solely on modeling network structures but fail to capture intrinsic features related to node importance, resulting in weak generalization across networks.
Our framework is a node-ranking method built on causal representation learning. Below, we review the two closely related research areas: node importance ranking and causal representation learning.
2.1 Node importance ranking
Previous works for identifying vital nodes in complex networks can be categorized into three main groups: traditional topological-based methods, machine learning-based methods and deep learning-based methods, as detailed in Table I. The first category, topology-based node importance estimation methods, mainly includes neighborhood-based, eigenvector-based, path-based and fusion-based approaches. Neighborhood-based methods, such as Degree Centrality (DC) [nieminen1974centrality], $H$-index [hirsch2005index], and $K$-shell decomposition [kitsak2010identification], assess node importance using local structural information. In contrast, eigenvector-based methods like Eigenvector Centrality (EC) [bonacich1987power] and PageRank [page1998pagerank] capture global influence propagation. Path-based methods, including Betweenness Centrality (BC) [freeman1977set], Eccentricity Centrality (ECC) [hage1995eccentricity], Closeness Centrality (CC) [freeman1978centrality] and ViralRank [iannelli2018influencers], evaluate nodes based on their positions and roles in network paths. While these methods offer computational simplicity and ease of interpretation, it is evident that these centrality measures depend solely on the network structure and capture only one specific structural aspect. Consequently, they still face challenges in scalability and adaptability when applied to more complex network structures. Fusion-based methods blend multiple centrality metrics or topological characteristics to create hybrid measures. Specifically, Mohammad et al. [mohammad2024convex] proposed a novel centrality measure called ISCL (Isolating Clustering Centrality), which employs a convex combination of isolating centrality and clustering coefficient with a tuning parameter to identify influential nodes in large-scale networks efficiently. Wang et al. [wang2024identification] introduced the HCT method, which combines a hypergraph fuzzy gravity model (HFGM), node centrality distribution characteristics (CDLW), and TOPSIS to progressively identify important nodes in multi-layer hypergraphs from local to global levels. Liu et al. [liu2023influence] proposed a Community-based Backward Generating Network (CBGN) framework for identifying influential nodes by integrating community detection, graph traversal, and an improved reverse-generation strategy to efficiently select high-impact seeds while minimizing influence overlap. An et al. [an2025identifying] proposed the Intrinsic and Topological Features-based Relationship Strength Gravity Centrality (ITFRSGC) that identifies key nodes in unweighted multi-layer networks by modeling hidden relationship strengths through multi-feature fusion and incorporating a dynamic influence range into a gravity centrality model. Nevertheless, these methods suffer from high computational complexity and sensitivity to parameter tuning, often requiring manual adjustment.
Machine learning–based approaches can partially overcome these issues by transforming the task into a regression or classification problem, using models like Linear Regression (LR), Multilayer Perceptron (MLP) and Support Vector Machine (SVM). Hajarathaiah et al. [hajarathaiah2024node] proposed a machine learning framework that combines traditional and local relative-change centrality metrics with epidemic simulation outcomes to estimate node importance. Wen et al. [wen2018fast] developed a two-level node ranking framework that integrates AHP-based evaluation with Least Squares Support Vector Machine (LS-SVM) learning to approximate global centrality measures using simple local indicators, enabling efficient and accurate node importance estimation. Wang et al. [wang2024multi] introduced a Multi-Factor Information Matrix (MFIM) that integrates node self-influence, neighbor influence, and mutual interaction through hybrid weighting to identify super-propagators in directed weighted social networks. While these methods significantly enhance computational efficiency and provide robust solutions, they still rely on explicit network topology information, which may limit their generalization performance in more complex scenarios.
Recent advancements in deep learning, particularly Graph Neural Networks (GNNs), have demonstrated superior performance by learning rich and informative representations of nodes. Zhao et al. [zhao2020infgcn] proposed the InfGCN algorithm that takes neighbor graphs and classic structural features as input into a graph convolutional network for learning node representations, followed by a classification layer. Zhang et al. [zhang2022new] combined convolutional neural networks (CNN) with GNNs in their CGNN algorithm, where nodes are labeled via the SIR model, and a loss function is optimized to accurately identify critical nodes. Ahmad and Wang [ahmad2024neural] proposed a Neural Diffusion Model (NDM) that leverages recurrent neural networks (RNNs) for dynamically modeling temporal influence propagation across networks. Wu et al. [wu2024graph] proposed ReGCN, a graph convolutional network model that leverages regular equivalence-based similarity to construct node features and employs an adaptive eigenvector selection strategy for influential node identification. Although these methods have shown considerable impact, they are inherently limited by their reliance on handcrafted features, which often reduces their ability to generalize across diverse network structures and capture complex structural patterns.
To overcome these limitations, kaili et al. [kaili2025finding] proposed GCRA, a deep learning-based framework that integrates Adaptive Graph Contrastive Learning (AdaGCL) with Deep Reinforcement Learning (DRL) to identify critical nodes by adaptively augmenting graph views and utilizing a dual-optimization reward function for network disintegration. Ma et al. [ma2025node] proposed a multi-view graph prompting framework (MVPL) that integrates node-level soft prompts and graph-level structural prompts to bridge the gap between pre-training and downstream node importance estimation tasks. Yu et al. [yu2020identifying] developed RCNN, a method that integrates GCNs with CNN-based adjacency features for efficient critical node detection. Xiong et al. [xiong2024vital] introduced the AGNN algorithm, which integrates autoencoders with GNNs to generate node embeddings using graph convolutional networks and optimizes ranking predictions based on the learned representations. Building on this work, Huang et al. [huang2025sparse] proposed a Sparse Transformer-enhanced Graph Convolutional Network (STGCN) to overcome AGNN’s limitations in large-scale and heterogeneous networks. STGCN integrates local topological modeling with global structural dependencies via stochastic anchor attention and masked mechanisms, enabling robust node-importance ranking in complex networks. Although these approaches bypass manual feature engineering by employing GNNs to learn topological embeddings and are capable of being pretrained on synthetic networks for ranking prediction on other real-world networks, they still face significant limitations. The learned representations are often produced by black-box models that lack an explicit and intrinsic connection to node importance, thereby limiting their generalizability across diverse networks. Our framework follows a similar paradigm to AGNN [xiong2024vital] and STGCN [huang2025sparse], in which models are trained on generative networks to enable ranking prediction on real-world networks. However, unlike these approaches, ICAN introduces a causal node embedding mechanism that can capture the causal relevance of representations to node importance, thereby enhancing both interpretability and generalization capability.
2.2 Causal representation learning
Causal relationships have the ability to capture the fundamental mechanism of data generation and are stable across various contexts [jiang2025integrating, zhou2025vulnerability]. Learning causal representations is significant for enhancing the robustness of predictive models. Zheng et al. [zheng2018dags] introduced NOTEARS, a novel method that reformulates the DAG discovery problem as a continuous optimization task incorporating acyclicity constraints, enabling solution via numerical techniques. Subsequently, Ng et al. [ng2019graph] expanded upon NOTEARS within a graph autoencoder framework, leveraging multilayer perceptrons (MLPs) to capture and model nonlinear structural equations, thereby advancing the capability to handle complex, non-linear relationships in causal inference. Yang et al. [yang2021learning] integrated a deep autoencoder and a causal structure learning model to learn causal representations using data. Yu et al. [yu2019dag] proposed DAG-GNN, a deep generative model that uses a graph neural network-based variational autoencoder to learn DAG structures from data, effectively generalizing linear structural equation models to capture nonlinear relationships and handle diverse variable types.
3 Problem formulation
@@
In this section, we define the problem of node importance ranking in complex networks and introduce the key assumptions of the proposed model.
3.1 Problem statement
@@
Given a graph $G=\{V,E\}$, where $V$ is the set of nodes and $E$ is the set of edges. The notation $v_{i}\in V$ denotes a node in $V$, and $e_{ij}\in E$ denotes an edge from node $v_{i}$ to $v_{j}$. The graph can be represented by an $n\times n$ adjacency matrix $\bm{A}$, where $n$ denotes the number of nodes. The entries of the matrix are defined such that $\bm{A}_{ij}=1$ when $e_{ij}\in E$, and $\bm{A}_{ij}=0$ otherwise. Let $\bm{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix. For graphs without node attributes, $\bm{X}$ can be obtained by applying a node embedding method (e.g., Node2Vec [grover2016node2vec]) to the adjacency matrix $\bm{A}$.
@@
The node importance ranking is a critical problem in network analysis. Given the graph $G=\{V,E\}$, our task is to assign a score $s_{i}$ to each node $v_{i}\in V$, forming a score vector $\bm{S}=(s_{1},s_{2},\dots,s_{n})$, These scores are then used to derive a node importance ranking $\bm{R}=(r_{1},r_{2},\dots,r_{n})$, where $r_{i}$ denotes the rank of node $v_{i}$. The ranking $\bm{R}$ satisfies the following: $r_{i}=r_{j}$ if $s_{i}=s_{j}$, and $r_{i}<r_{j}$ if $s_{i}<s_{j}$, meaning that a higher score corresponds to a higher (i.e., numerically lower) rank. Unlike methods that emphasize the absolute magnitude of scores, our approach focuses on preserving the relative ordering of nodes consistent with their underlying importance, making ranking consistency the primary objective.
3.2 Assumptions